OpenStudy (anonymous):

dr/ d'theta theta=3 r=4/squareroot(12-theta)

OpenStudy (anonymous):

\[r=(4)(12+\theta)^-1/2\]4(-1/2)[12-\theta]^[-1/2-1] \[dr/d \theta=2[12-\theta]^-3/2\] when \theta=3 \[dr/d \theta =2/[12-3]^3/2=2/(9)\sqrt{9}=2/27\]

OpenStudy (anonymous):

\[y=\cos(x) -2\pi \le x \le2\pi x=3pi/2]